# Network-Realized Model Predictive Control Part II: Distributed Constraint Management

### v1

@@ Abstract
A two-layer control architecture is proposed, which promotes scalable implementations for model predictive controllers. The top layer acts as both reference governor for the bottom layer, and as a feedback controller for the regulated network. By employing set-based methods, global theoretical guarantees are obtained by enforcing local constraints upon the variables of the network and of the first layer’s implementation. The proposed technique offers recursive feasibility guarantees as one of its central features, and the expressions of the resulting predictive strategies bear a striking resemblance to classical formulations from model predictive control literature, allowing for flexible and easily customizable implementations.
@@
Distributed control; convex optimization; model predictive control; networked systems; scalable implementations.
@@ 1 Introduction
1.1 Context and Motivation
Set-based constraint satisfaction is one of the key features of Model Predictive Control (MPC, see [3], [7] and [4]). Given the steady increase in popularity of these predictive strategies, recent results in system-theoretical literature have focused on expanding these techniques to a distributed setting, in which sensor information and access to actuators is only partially available in any given location of the controlled network. In spite of being readily apparent, the possibility of implementing separate MPC-based subcontrollers is met with a daunting challenge: since the command signals are computed independently of one another, they do not inherently take into account the effect of the cross-coupling induced by their control action. This fact prompted the development of more sophisticated procedures, such as the ones from [1] and [2], in which the aforementioned cross-coupling is attenuated by a prominent parametrization of distributed control laws (see [12]). However, since the aforementioned controllers are synthesized while taking into account MPC-like cost functions and constraints, any modification to the latter leads to their implicit redesign.
@@ 1.2 Scope of Work
The problem being tackled in this manuscript, and in its companion paper [10], is that of employing a class of distributed controllers (see [13] for a general overview) to facilitate the design of scalable and computationally efficient MPC-based policies for networks of widely distributed systems. By building upon the closed-loop dynamics obtained using the Network Realization Functions (NRF) formalized in [9] and [11], the problem of obtaining area-based prediction models is addressed in the context of distributed synthesis for MPC-based policies. The proposed design framework involves using these dynamically decoupled and stable models in order to solve local subproblems, with a focus on constraint handling, while addressing objectives that pertain only to an area’s assigned variables. By leveraging a number of set-theoretical results (as treated in [3]), and by reinterpreting them to our distributed setting, the resulting control actions are meant to offer global guarantees when applied in tandem to the closed-loop system formed by the network and the first layer.
1.3 Contributions
This paper focuses on the MPC layer of our control architecture, while also incorporating the structured dynamics obtained via the NRF-based implementation of the first layer. Consequently, we present a novel perspective on MPC-based design, by exploiting the ability of these control policies to act simultaneously as both reference governor [5] for the first layer’s subcontrollers and as regulatory strategies for the controlled network’s original dynamics. In effect, we propose computationally efficient means of influencing a closed-loop system’s evolution, in order to ensure the satisfaction of constraints for both the network’s actuation signals and for its state variables. We achieve this in a distributed setting, in which we extend the theoretical framework of [6] to an MPC-like setting: one in which ensuring local constraint satisfaction, using only proximally available information and means of actuation, leads to the fulfilment of global constraints. Finally, we show that even the recursive feasibility of our solution can be successfully obtained by employing the same local-to-global theoretical framework.
@@ 1.4 Paper Structure
The rest of the paper is structured as follows. Section 2 contains a set of preliminary notions, while Section 3 discusses our two-layer architecture and its overarching objectives, with a particular focus on the MPC layer. Section 4 presents the means of obtaining tractable prediction models for a area-partitioned control scheme, and Section 5 handles the technicalities of constraint management in the distributed setting. Section 6 holds the main theoretical results and implementation procedures, while Section 7 showcases a numerical example based upon the vehicle platooning application discussed in [8] and Section 8 offers a set of concluding remarks.
@@ 2.1 Notation and Definitions
Let $\mathbb{N}$, $\mathbb{R}$ and $\mathbb{C}$ denote, respectively, the set of natural, real and complex numbers. Let $\mathcal{M}^{p\times m}$ be the set of all $p\times m$ matrices whose entries are scalar elements that part of a set denoted by $\mathcal{M}$. Similarly, $\mathcal{M}^{p}$ denotes the set of vectors with dimension $m$ and entries in $\mathcal{M}$. For any matrix $M$, $\mathrm{row}_{i}(M)$ is its $i^{\text{th}}$ row, $\mathrm{col}_{j}(M)$ stands for its $j^{\text{th}}$ column and $\mathrm{elm}_{ij}(M):=\mathrm{row}_{i}(\mathrm{col}_{j}(M))$. Moreover, the transpose of an arbitrary matrix $M$ is denoted by $M^{\top}$. We denote by $\mathcal{R}(z)$ the set of real-rational functions of indeterminate $z\in\mathbb{C}$. A matrix with entries in $\mathcal{R}(z)$ is termed a Transfer Function Matrix (TFM), and all such TFMs will be denoted using boldface letters.
The vector $e_{i}$ stands for the $i^{\text{th}}$ term in the canonical basis of $\mathbb{R}^{m}$, for some $\0m\in\mathbb{N}\backslash\{0\}$. For any $M\in\mathbb{R}^{p\times m}$ and any $\mathcal{X}\subseteq\mathbb{R}^{m}$, $M\mathcal{X}\subseteq\mathbb{R}^{p}$ denotes the image of $\mathcal{X}$ under $M$ and $(-\mathcal{X}):=\{-x\ |\ x\in\mathcal{X}\}$. Moreover, for any two subsets $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ of the same set $\mathcal{X}\subseteq\mathbb{R}^{m}$, $\mathcal{X}_{1}\oplus\mathcal{X}_{2}$ represents the Minkowski sum of the subsets, while $\mathcal{X}_{1}\ominus\mathcal{X}_{2}$ denotes the Pontryagin difference of $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ (see, for example, Definition 3.10 in [7]).
@@ 2.2 System Representations
for any $k\in\mathbb{N}$ with $k\geq k_{0}\in\mathbb{N}$, where $k_{0}$ represents the initial time instant. For the representations of type (1a)-(1b), which are commonly referred to as state-space realizations, $u$ is the realization’s controlled input vector, $d$ its uncontrolled input vector or, alternatively, its disturbance vector, $y$ its output vector and $x$ its state vector.
For such systems, we proceed to denote by $x_{c}\in\mathbb{R}^{n_{x}}$ the initial condition of the state vector, and we also consider $A\in\mathbb{R}^{n_{x}\times n_{x}}$, $B_{u}\in\mathbb{R}^{n_{x}\times n_{u}}$, $B_{d}\in\mathbb{R}^{n_{x}\times n_{d}}$, $C\in\mathbb{R}^{n_{y}\times n_{x}}$, while $D_{u}\in\mathbb{R}^{n_{y}\times n_{u}}$ and $D_{d}\in\mathbb{R}^{n_{y}\times n_{d}}$. The positive integer $n_{x}$ in (1a)-(1b) is called the order of the realization and we denote any such realization in compact form via a matrix quadruplet of type $\left(A,\begin{bmatrix}B_{u}&B_{d}\end{bmatrix},C,\begin{bmatrix}D_{u}&D_{d}\end{bmatrix}\right)$.
The link between a system which is described by (1a)-(1b) and its TFM is given by
$\mathbf{G}(z)=\left[\scriptsize\begin{array}[]{c|cc}A-zI_{n_{x}}&B_{u}&B_{d}\\ \hline\cr C&D_{u}&D_{d}\end{array}\right]:=\\ :=C(zI_{n_{x}}-A)^{-1}\scriptsize\begin{bmatrix}B_{u}&B_{d}\end{bmatrix}+\scriptsize\begin{bmatrix}D_{u}&D_{d}\end{bmatrix}.$ (2)
Finally, given any proper $\mathbf{H}\in\mathcal{R}(z)^{n_{y}\times n_{u}}$, the notation $y[k]=\mathbf{H}(z)\star u[k]$ denotes the time-response of $\mathbf{H}(z)$ to an input signal vector $u[k]$, which can be computed as
@@
for any realization $(A_{H},B_{H},C_{H},D_{H})$ of $\mathbf{H}(z)$.
[…]
@@ 2.3 Distributed Networks
The type of networks our procedure is catered towards are described by a more particular class of realizations than the ones from (1a)-(1b), namely
$x[k+1]=$ $\ Ax[k]+B_{u}u[k]+B_{d}\,d[k],$ (4a)
$y[k]=$ $\ \phantom{A}x[k].$ (4b)
Although we consider that $x$ and $u$ are completely accessible to us, in terms of measurement and actuation respectively, we do not assume that manipulating all of this variables is possible from any one location in our physical network. This naturally separates the network into a number of $N\in\mathbb{N}$ distinct areas, with $N>1$.
To each of the network’s $N$ areas we now assign two sets of indices, which denote the entries of $x$ and $u$ that are (uniquely) accessible for measurement or actuation in that area. Without loss of generality, we assume that these indices are assigned to the sets in ascending order, since the rows and columns of the matrices expressed in (4a)-(4b) may be permuted at will. Therefore, each area will be represented by the pair
$\mathcal{A}_{i}:=(\mathcal{A}_{xi},\mathcal{A}_{ui}),\,\forall\,i\in\{1:N\},\vspace{-2mm}$ (5)
@@
$\left\{\begin{split}\mathcal{A}_{\bullet i}:=\{\alpha_{\bullet i}+j\ |\ j\in 1:n_{\bullet i}\},\,\forall\,i\in\{1:N\},\\ \alpha_{\bullet 1}:=0,\,\alpha_{\bullet\ell}:=\alpha_{\bullet(\ell-1)}+n_{\bullet(\ell-1)},\,\forall\,\ell\in\{2:N\}.\end{split}\right.\vspace{-2mm}$ (6)
To each partition [^1] As an example, a network with 12 states and 7 inputs could be partitioned as in (5)-(6) via the following collection of sets {($\scriptsize\begin{array}[]{l}\{(\{1:2\},\{1\}),(\{3:5\},\{2:3\}),(\{6:11\},\{4:5\}),(\{12\},\{6:7\})\}\end{array}$ in (6), we attach the selection matrix
$S_{\bullet i}:=\left[\scriptsize\begin{array}[]{ccc}e_{\alpha_{\bullet i}+1}&\dots&e_{\alpha_{\bullet i}+n_{\bullet i}}\end{array}\right],\,\forall\,i\in\{1:N\},\vspace{-2mm}$ (7)
@@ 3 The Control Architecture
3.1 Global Perspective and Objectives
The aim of this paper and of its companion paper [10] is to formalize the design framework of the two-layer control architecture for the interconnected dynamical system depicted in Figure 1. By identifying the subset of areas $\mathcal{N}_{i}\subseteq\{1:N\}$ which are able to transmit information to the network’s $i^{\text{th}}$ area (and which trivially satisfy $i\in\mathcal{N}_{i}$), for all $i\in\{1:N\}$, the communication graph is clearly designated and the fundamental operating principle of our control scheme is the following:
@@ 1. $a)$
    the state-space implemented NRF subcontrollers receive network state information along with the command signals of other first layer subcontrollers from their areas’ neighbourhoods, and use these values to compute the signals denoted $u_{fi}$ in Figure 1;
@@ 2. $b2)$
In terms of control architecture design, it should be clear that both the first and the second layer employ feedback and feedforward components. However, their goals and their communication-based mechanisms are completely different. Where the first layer focuses on dynamical decoupling and disturbance rejection, the second layer concerns itself with constraint satisfaction and feasibility, in terms of robust control invariance.
The architecture and design of the first layer is covered in our companion paper [10], and this manuscript focuses on all aspects concerning the MPC-like second layer of our proposed control scheme. However, given that the MPC layer is built upon the closed-loop interconnection between the NRF layer and the network, its objectives will invariably account for the dynamics of the NRF-based subcontrollers. Thus, we briefly summarize several key concepts from [10], before laying out the main objectives of our top-layer predictive strategies.

### v4

# Network-Realised Model Predictive Control Part II: Distributed Constraint Management

@@ Abstract
A two-layer control architecture is proposed, which promotes scalable implementations for model predictive controllers. The top layer acts as both reference governor for the bottom layer, and as a feedback controller for the regulated network. By employing set-based methods, global theoretical guarantees are obtained by enforcing local constraints upon the network’s variables and upon those of the first layer’s implementation. The proposed technique offers recursive feasibility guarantees as one of its central features, and the expressions of the resulting predictive strategies bear a striking resemblance to classical formulations from model predictive control literature, allowing for flexible and easily customisable implementations.
@@
Distributed control; convex optimisation; model predictive control; networked systems; scalable implementations.
@@ 1 Introduction
1.1 Setting and Context
Model Predictive Control (MPC, see [RMD] and also [BBM]) is one of the most popular control-based frameworks in both the industrial setting and in system-theoretical literature. One of the key features that has generated its current popularity is its affinity for enforcing set-based constraints in the centralised setting [STMC]. On the other hand, as technical applications are becoming increasingly interconnected and widely distributed, the synthesis of distributed control laws has evolved into one of the foci of contemporary research for the control community.
Among the myriad approaches developed for the MPC paradigm (see [DMPCB] for a comprehensive survey on the topic), some of the most promising solutions are those inspired by robust MPC (see [RMPC] for a popular centralised setting), among which we highlight the techniques proposed in [R8A, R8B, R8C, R8D] for the distributed setting. A key aspect of this particular framework is given by the use of an ancillary control law, based upon static state feedback, which operates in tandem with its MPC policy.
Distinct from the latter approach, the procedure formalised in [DMPC1, DMPC2] aims to leverage the powerful design framework from [SLS] in order to tackle the intricacies of the distributed setting. One such intricacy, which arises chiefly in formation control, is the problem of attenuating the propagation of disturbance through said formation (see [SS_rev]). This problem has received a great deal of attention in MPC-oriented literature, with multiple solutions having been proposed in both the distributed setting [plat_MPC_1, plat_MPC_2, plat_MPC_3, plat_MPC_4], and also in the centralised one [plat_MPC_5].
1.2 Motivation
Despite the valuable contributions made in the previously cited works, certain limitations which are inherent in these approaches serve as the chief motivation for the developments presented in this paper. More precisely, the control laws proposed in [R8A, R8B, R8D] notably require a set of static and stabilising state-feedback strategies, which are constrained to having block-diagonal sparsity structures. Furthermore, the procedure described in [R8C] revolves around intricate set-theoretical notions, such as joint set invariance, while implicitly assuming that the controlled network is bipartite and that the state matrix of its (linear) dynamics satisfies a restrictive spectral condition. Finally, the design framework presented in [DMPC1, DMPC2] tackles both dynamical disturbance decoupling and distributed constraint satisfaction jointly, in a single design phase, whereas the formation control-oriented approaches in [plat_MPC_1, plat_MPC_2, plat_MPC_3, plat_MPC_4, plat_MPC_5] require restrictive or computationally demanding procedures to reach their goals.
@@ 1.3 Scope of Work
In contrast to the techniques from [R8A, R8B, R8C, R8D], the MPC-based control laws discussed in this paper do not focus on network regulation in the presence of constraints, but rather on closed-loop supervision of both the plant and its control laws, which based upon the Network Realisation Function (NRF, see the companion paper [part1], along with [NRF]) representation. In our framework, disturbance rejection and reference tracking are handled predominantly by the distributed (and dynamical) controllers discussed in [part1]. By exploiting the closed-loop properties ensured by these NRF-based subcontrollers [plutonizare, NRF, aug_sparse], the MPC-inspired procedures stated in the sequel are more akin to distributed versions of the so-called reference governors (see [refgov] for a broad discussion on these policies). Indeed, a careful inspection of the survey presented in [arch_rev] reveals that our framework does not fit neatly into any of the architectures addressed therein.
1.4 Contribution
From a system-theoretical perspective, the chief contributions of the work presented in this paper can be listed via the following three features:
  1. a)
    ensuring command constraint satisfaction, in an MPC-based framework, for a plant in closed-loop configuration with a dynamical controller;
  2. b)
    coordinating the control authority of our MPC-based policies with the NRF-based controllers from [NRF, aug_sparse], while splitting the MPC’s actuation efforts between the latter and the controlled network;
  3. c)
    enabling constraint-compliant and NRF-based control, in which the MPC-based subcontrollers are designed to act only as a safeguard mechanism against constraint violation.
A further point to emphasise is the fact that all of the above features are achieved in a fully distributed setting, and we point out that the mechanism by which a dynamical controller’s output is reliably constrained (for the purpose of command limitation) has not, to the best of our knowledge, been adequately addressed in system-theoretical literature, even in the centralised setting.
@@ 1.5 Paper Structure
The rest of the paper is structured as follows. Section 2 contains a set of preliminary notions, while Section 3 discusses our two-layer architecture and its overarching objectives, focusing on the MPC layer. Section 4 presents the means of obtaining tractable prediction models for an area-partitioned control scheme, and Section 5 handles the technicalities of constraint management in the distributed setting. Section 6 holds the main theoretical results and implementation procedures, while Section 7 showcases a numerical example based upon the vehicle platooning application discussed in [plutonizare] and Section 8 offers a set of concluding remarks. In addition to this, we include a pair of appendices: one focusing on the key concepts of NRF-based control laws (Appendix A), as discussed in our companion paper [part1], and one which holds the proofs of this paper’s main results (Appendix B).
We now move on to introducing the various preliminary notions which are necessary to state our main results.
@@ 2.1 Notation and Definitions
Let $\mathbb{N}$, $\mathbb{R}$ and $\mathbb{C}$ denote, respectively, the set of natural, real and complex numbers. Let $\mathcal{M}^{p\times m}$ be the set of all $p\times m$ matrices whose entries are scalar elements that are part of a set denoted $\mathcal{M}$. Similarly, $\mathcal{M}^{p}$ denotes the set of vectors with dimension $p$ and entries in $\mathcal{M}$. For any matrix $M$, $\mathrm{row}_{i}(M)$ is its $i^{\text{th}}$ row, $\mathrm{col}_{j}(M)$ stands for its $j^{\text{th}}$ column and $\mathrm{elm}_{ij}(M):=\mathrm{row}_{i}(\mathrm{col}_{j}(M))$. Moreover, the transpose of an arbitrary matrix $M$ is denoted $M^{\top}$. We denote by $\mathcal{R}(z)$ the set of real-rational functions of indeterminate $z\in\mathbb{C}$. A matrix with entries in $\mathcal{R}(z)$ is called a Transfer Function Matrix (TFM), and all such TFMs will be denoted using boldface letters.
The vector $e_{i}$ stands for the $i^{\text{th}}$ column belonging to the identity matrix of compatible dimensions. For any matrix $M\in\mathbb{R}^{p\times m}$ and any set $\mathcal{X}\subseteq\mathbb{R}^{m}$, we define $M\mathcal{X}:=\{Mx\ |\ x\in\mathcal{X}\}$ and $(-\mathcal{X}):=\{-x\ |\ x\in\mathcal{X}\}$. In addition to this, for any two sets $\mathcal{X}_{1}\subseteq\mathbb{R}^{m}$ and $\mathcal{X}_{2}\subseteq\mathbb{R}^{m}$, we will henceforth denote their Minkowski sum as follows $\mathcal{X}_{1}\oplus\mathcal{X}_{2}:=\{x_{1}+x_{2}\ |\ x_{1}\in\mathcal{X}_{1},\ x_{2}\in\mathcal{X}_{2}\}$, while the operation $\mathcal{X}_{1}\ominus\mathcal{X}_{2}:=\{y\in\mathbb{R}^{m}\ |\ \{y\}\oplus\mathcal{X}_{2}\subseteq\mathcal{X}_{1}\}$ will represent the Pontryagin difference between the two sets.
@@ 2.2 System Representations
for any $k\in\mathbb{N}$ with $k\geq k_{0}\in\mathbb{N}$, where $k_{0}$ represents the initial time instant. For the representations of type (1a)-(1b), which are referred to as state-space realisations, $u$ is the realisation’s controlled input vector, $d$ its disturbance vector, $y$ its output vector and $x$ its state vector.
The type of network that our procedure aims to control is described by a more particular class of realisations than (1a)-(1b). These systems also satisfy the fact that
$C=I,\ D_{u}=O,\ D_{d}=O,\vskip-8.53581pt$ (1c)
implying that the states are available [^1] While being subject to measurement noise, as will be discussed (and suitably compensated for) in the sequel. for measurement.
Remark 2.1.
The disturbance vector $d[k]$ along with the noise affecting the state measurements (1b)-(1c) are key components of our framework. One of the main goals of the distributed MPC-based procedure proposed in this paper is to ensure constraint satisfaction for $x[k]$ and $u[k]$ in the presence of the aforementioned exogenous signals (along with additional ones, introduced in the sequel).
For systems of type (1a)-(1c), we denote by $x_{c}\in\mathbb{R}^{n_{x}}$ the initial condition of the state vector and we also consider $A\in\mathbb{R}^{n_{x}\times n_{x}}$, $B_{u}\in\mathbb{R}^{n_{x}\times n_{u}}$, $B_{d}\in\mathbb{R}^{n_{x}\times n_{d}}$, $C\in\mathbb{R}^{n_{y}\times n_{x}}$, while, more generically, we consider $D_{u}\in\mathbb{R}^{n_{y}\times n_{u}}$ and $D_{d}\in\mathbb{R}^{n_{y}\times n_{d}}$. The positive integer $n_{x}$ will be referred to as the order of the realisation, and the link between a system described by (1a)-(1b) and its TFM is given by
$\mathbf{G}(z)=\left[\scriptsize\begin{array}[]{c|cc}A-zI_{n_{x}}&B_{u}&B_{d}\\ \hline\cr C&D_{u}&D_{d}\end{array}\right]:=\\ :={\scriptsize\begin{array}[]{l}C(zI_{n_{x}}-A)^{-1}\end{array}}\hskip-2.84526pt\scriptsize\begin{bmatrix}B_{u}&B_{d}\end{bmatrix}+\scriptsize\begin{bmatrix}D_{u}&D_{d}\end{bmatrix}$. (2)
Finally, given any $\mathbf{H}\in\mathcal{R}(z)^{n_{y}\times n_{u}}$ which can be described by (1a)-(1b) and (2), the shorthand notation $y[k]=\mathbf{H}(z)\star u[k]$ represents the time-response of $\mathbf{H}(z)$ to an input signal vector $u[k]$, which can be computed as
@@
for any realisation $(A_{H},B_{H},C_{H},D_{H})$ of $\mathbf{H}(z)$.
InterconnectedDynamicalSystemNRF 11Area $\mathbf{1}$MPC $1x_{1}$$u_{f1}$$\{x_{1},w_{1}\}$$u_{s11}$$u_{s21}$$u_{f1}$$\{x_{j\in\mathcal{N}_{1},j\neq 1},u_{f(j\in\mathcal{N}_{1},j\neq 1)},u_{s1(j\in\mathcal{N}_{1},j\neq 1)}\}$$\{x_{1},w_{1},u_{s11}\}$$\{x_{j\in\mathcal{N}_{1},j\neq 1},w_{j\in\mathcal{N}_{1},j\neq 1}\}\cdotsNRF$ iiArea $\mathbf{i}$MPC $ix_{i}$$u_{fi}$$\{x_{i},w_{i}\}$$u_{s1i}$$u_{s2i}$$u_{fi}$$\{x_{j\in\mathcal{N}_{i},j\neq i},u_{f(j\in\mathcal{N}_{i},j\neq i)},u_{s1(j\in\mathcal{N}_{i},j\neq i)}\}$$\{x_{i},w_{i},u_{s1i}\}$$\{x_{j\in\mathcal{N}_{i},j\neq i},w_{j\in\mathcal{N}_{i},j\neq i}\}\cdots\cdots$
@@ 2.3 Distributed Networks
Although we consider that the vectors $x$ and $u$ are completely accessible, in terms of measurement and actuation, respectively, we do not assume that manipulating all of these variables is possible from any one location in our physical network. This naturally separates the network into a number of $N\in\mathbb{N}$ distinct areas, with $N>1$.
To each of the network’s $N$ areas, we now assign two sets of indices, which denote the entries of $x$ and $u$ that are (uniquely) accessible for measurement or actuation in that area. Without loss of generality, we assume that these indices are assigned to the sets in ascending order, since the rows and columns of the matrices expressed in (1a)-(1c) may be permuted at will. Therefore, each area will be represented by the pair
$\mathcal{A}_{i}:=(\mathcal{A}_{xi},\mathcal{A}_{ui}),\,\forall\,i\in\{1:N\},\vskip-8.53581pt$ (4)
@@
[…]
To each partition [^2] As an example, a network with 12 states and 7 inputs could be partitioned as in (4)-(5) via the following collection of sets $\scriptsize\begin{array}[]{l}\{(\{1:2\},\{1\}),(\{3:5\},\{2:3\}),(\{6:11\},\{4:5\}),(\{12\},\{6:7\})\}\end{array}$ in (5), we attach the selection matrix
$S_{\bullet i}:=\left[\scriptsize\begin{array}[]{ccc}e_{\alpha_{\bullet i}+1}&\dots&e_{\alpha_{\bullet i}+n_{\bullet i}}\end{array}\right],\,\forall\,i\in\{1:N\},\vskip-5.69054pt$ (6)
@@ 3 The Control Architecture
3.1 Global Perspective and Functioning
The aim of this manuscript and of its companion paper [part1] is to formalise the design framework of the two-layer control architecture for the interconnected dynamical system depicted in Figure 1. By identifying the subset of areas $\mathcal{N}_{i}\subseteq\{1:N\}$ which are able to transmit information to the network’s $i^{\text{th}}$ area (and which trivially satisfy $i\in\mathcal{N}_{i}$), for all $i\in\{1:N\}$, the communication graph is clearly designated and the fundamental operating principle of our control scheme is the following:
@@ 3.1 Global Perspective and Functioning
    the state-space-implemented and NRF-based subcontrollers receive network state feedback and the command signals of other first-layer subcontrollers from their areas’ neighbourhoods, using these values to compute the signals denoted $u_{fi}$ in Figure 1;
@@
In terms of control architecture design, it should be clear that both the first and the second layers employ feedback and feedforward components. However, their goals and their communication-based mechanisms are completely different. Where the first layer focuses on dynamical decoupling and disturbance rejection, the second layer concerns itself with constraint satisfaction and feasibility, in the sense of robust control invariance.
As previously mentioned in Remark 2.1, disturbance plays a key role in our chosen framework. Thus, we recall the disturbance vector $d[k]$ from (1a)-(1c) and we denote the measurement noise mentioned in said remark as $\zeta[k]\in\mathbb{R}^{n_{x}}$. Moreover, we also introduce $\beta_{f}[k]\in\mathbb{R}^{n_{u}}$, for the first layer’s communication disturbance, along with $\beta_{s1}[k]\in\mathbb{R}^{n_{x}}$ and $\beta_{s2}[k]\in\mathbb{R}^{n_{u}}$, the communication disturbance associated with the second layer. The latter three signal vectors act as additive disturbance on the command signals depicted in Figure 1, namely $u_{f}[k]:=\tiny\begin{bmatrix}u_{f1}^{\top}[k]&\dots&u_{fN}^{\top}[k]\end{bmatrix}^{\top}$, $u_{s1}[k]:=\tiny\begin{bmatrix}u_{s11}^{\top}[k]&\dots&u_{s1N}^{\top}[k]\end{bmatrix}^{\top}$ and $u_{s2}[k]:=\tiny\begin{bmatrix}u_{s21}^{\top}[k]&\dots&u_{s2N}^{\top}[k]\end{bmatrix}^{\top}$, respectively. All of the exogenous signals mentioned in this paragraph may be combined into the global disturbance vector
[…]
with respect to which we now proceed to formally define the objectives of our MPC layer.
3.2 Objectives of the MPC Layer
The second layer of the control architecture described in Section 3.1 must be formed by $N$ separate MPC-based subcontrollers, which accomplish the following goals:
  1. I)
    Given two sets $\mathcal{X}\subseteq\mathbb{R}^{n_{x}}$ and $\mathcal{U}\subseteq\mathbb{R}^{n_{u}}$, the second layer’s global control signals $u_{s1}[k]$ and $u_{s2}[k]$ robustly satisfy the constraints $x[k]\in\mathcal{X}$ and $u[k]\in\mathcal{U}$ in the presence of $d_{s}[k]$ from (7);
  2. II)
    The area-based components of $u_{s1}[k]$ and $u_{s2}[k]$ are computed using receding horizon-based optimisation (see, for example, Chapter 1 in [RMD]), with each area admitting tractable formulations of type (8), located at the top of this page, where:
    1. a)
      $g_{t}(\cdot)$ is the stage cost at time instant $k+t$,
    2. b)
      $\xi_{i}$ denotes the local state vector of the $i^{\text{th}}$ area’s closed-loop system, to be addressed (along with its particular initialisation) in the sequel,
    3. c)
      the scalar values $T_{i}\in\mathbb{N}\setminus\{0\}$ and $\overline{T}_{i}\in\mathbb{N}$ are, respectively, the (set-)constrained and the unconstrained prediction horizons of each area,
    4. d)
      the sets denoted ${\mathcal{U}}_{s1i}$ along with ${\mathcal{U}}_{s2i}$ constrain the local second-layer commands and $\Xi_{it}$ is the set linking the local constraints with the global ones, to be constructed in the sequel;
  3. III)
    For all $i\in\{1:N\}$, the $i^{\text{th}}$ area’s MPC-based subcontroller from (8) computes $u_{s1i}[k]$ and $u_{s2i}[k]$ using only information available in $\mathcal{N}_{i}$;
  4. IV)
    Ensuring the feasibility of each optimisation problem from (8) at some $k=k_{0}$ ensures all of them remain feasible for all $k>k_{0}$.
Remark 3.2.
As previously indicated in Remark 3.1, the goals stated above show that our second layer is only charged with enforcing constraints in a distributed manner. Thus, all objectives pertaining to reference tracking or robust stabilisation are to be handled by the NRF-based first layer, in order to take full advantage of its remarkable closed-loop guarantees (see, for example, the main results of [plutonizare, NRF, aug_sparse]). In our framework, the MPC layer acts primarily as a safeguard for the NRF-based control laws, further highlighting the particular nature of our approach. As shown by the numerical example in the sequel, we aim to design our second layer so as to avoid acting on the first-layer-based closed-loop system as much as possible, only generating command signals of non-negligible amplitude when constraint violation becomes imminent.
In order to achieve all of the above-stated objectives, however, tractable and area-based models will have to be derived for the first-layer-based closed-loop system.
4 Obtaining an MPC-Oriented Formulation
This section is dedicated to discussing the derivation and the manipulation of the models which form the equality constraints of the problems stated in (8). Readers familiar with the robust MPC framework (see Chapter 3 in [RMD] for a comprehensive treatment of the topic) will undoubtedly recognise many of the formulations and conventions adopted in the sequel. Nevertheless, the usage of dynamical feedback instead of the classical, static gains in the so-called tube-based MPC framework (see, for example, Section 3.5 in [RMD]) renders our problem distinct from other approaches in system-theoretical literature. Therefore, given this distinct element of novelty, we proceed to lay out our expressions in an explicit manner.
4.1 Interfacing with the First Layer
We begin by briefly summarising several key concepts from our companion paper [part1]. Our architecture’s first layer is implemented as depicted in Figure 2, located on the next page, in which the distributed control laws are represented by the TFM
[…]
where each $\mathbf{K}_{\mathbf{D}i}(z)=S_{ui}^{\top}\mathbf{K}_{\mathbf{D}}(z)$ represents the TFM of the $i^{\text{th}}$ area’s local subcontroller.
Remark 4.1.
We operate under the assumption that the communication infrastructure of our control architecture is endowed with error-detecting/correcting transmission mechanisms. Thus, all communication errors affecting transmitted information are modelled as disturbance which originates at the source of transmission, not at the receiving ends. Consequently, the resulting signal diagram of the first layer’s closed-loop system will be precisely the one shown in Fig. 2, even when all $\mathbf{K}_{\mathbf{D}i}(z)$ are implemented separately, and the closed-loop mappings will have the beneficial properties discussed in [NRF, part1, aug_sparse].
$\mathbf{K}_{\mathbf{D}i}(z)=\left[\scriptsize\begin{array}[]{c|c}A_{wi}-zI_{n_{wi}}&B_{wi}\\ \hline\cr C_{wi}&D_{wi}\end{array}\right]:=\left[\tiny\begin{array}[]{ccc|c}A_{r(\alpha_{ui}+1)}-zI_{n_{r(\alpha_{ui}+1)}}&&&B_{r(\alpha_{ui}+1)}\\ &\ddots&&\vdots\\ &&A_{r(\alpha_{ui}+n_{ui})}-zI_{n_{r(\alpha_{ui}+n_{ui})}}&B_{r(\alpha_{ui}+n_{ui})}\\ \hline\cr C_{r(\alpha_{ui}+1)}&&&O\\ &\ddots&&\vdots\\ &&C_{r(\alpha_{ui}+n_{ui})}&O\\ \end{array}\right].$ (10)
Crucially, these subcontrollers are implemented via the state-space realisations given in (10), located at the top of the next page, whose component matrices (see Section 4.1 in [part1] for more details) are given, $\forall\,\ell\in\mathcal{A}_{ui}$, by
[…]
Henceforth, we define the orders $n_{wi}:=\textstyle\sum_{\ell=1}^{n_{ui}}n_{r\ell}$ and $n_{w}:=\textstyle\sum_{i=1}^{N}n_{wi}$. We also denote the state vectors of the implementations from (10) as $w_{i}[k]$ and their initial conditions as $w_{ci}\in\mathbb{R}^{n_{wi}}$. Moreover, we concatenate the latter into $w_{c}:=\tiny\begin{bmatrix}w_{c1}^{\top}&\dots&w_{cN}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{n_{w}}$, which are the initial conditions of the first layer’s global implementation.